[Hallc_running] [Revised Logentry] RC Update - July 19, 2025
kinney at jlab.org
kinney at jlab.org
Sat Jul 19 09:25:02 EDT 2025
Logentry Text:
--
Major progress was made during the data as the coincidence time calculation from the ROC readouts (see log entry https://logbooks.jlab.org/entry/4420966). Thanks to Mark Jones for coming out to help us! Estimated yields are now much closer to those estimated by Peter and Provakar.
Bill Henry provided a temporary fix for the broken cable connection on one end of an SHMS hodoscope paddle (https://logbooks.jlab.org/entry/4421094). After discussion at the 2:30 meeting, we agreed with Bill that the FADC offsets needed adjustment, which he then implemented (https://logbooks.jlab.org/entry/4421084). I believe this means that reference time cuts need to be checked/changed, but time difference cuts should not need changing.
We had a semi-major flail when something caused all the large status screens to disappear. We fear that someone working on the machine was killing processes unaware that they were in use in the counting house. If you work remotely on those machines, please be careful!
We also had a hiccup in the pre-scale GUI which was failing to update CODA correctly (setting of prescale rate did not change coda event rate, but did change livetime dramatically). Solution suggested by Dave G. worked. Just stop run, kill and restart prescale GUI (go_prescale_SHMS) corrected the situation. It's a good idea to check that prescale rate and coda rate agree!
Data collection continued on pi- SIDIS data at x=0.25, Q2=3.3, z=0.36 through day, swing and owl shift. The accelerator turned off for a relatively brief time (< 1hr) during the passage of a thunderstorm overhead. During the owl shift, we completed measurements of DIS in the HMS and SHMS for all the targets, then returned to finish the remaining cycles of SIDIS at this kinematic. Beam has generally been very good, even with the higher currents used in the dummy runs.
After finishing the z=0.36 pi- SIDIS kinematic, we will move to the z=0.9 pi- SIDIS. At 2pm we hope to perform a BCM calibration with Dave Mack's help.
There will be an RC meeting today at 2:30 pm in the meeting room on the 2nd floor of the Counting House. The usual Zoom information is below:
https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu...
Meeting ID: 161 822 7115
Passcode: 253463
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4421733
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250719/98ed5447/attachment.htm>
More information about the Hallc_running
mailing list